Hertz car rental agency charges $25 daily plus $.10 per mile. Budet charges $20 daily plus $.20 per mile. Find the daily mileage for which the Budget charge is equal to the Hertz charge. (Please can u show me how to do this step by step using the substitution method?)
Answer by mananth(16949)

Hertz Car $25+0.10x where x is the number of miles.
Budget car $20+0.20x
Both equal
25+0.1x=20+0.2x
25-20 = 0.2x-0.1x
5=0.1x
50=x
50 miles